# thanks: Equal contributionsthanks: Equal contributions

### v1

@@ Abstract
Viscous films flowing down an incline can form self-sustained running waves, known as Kapitsa roll waves. Here we describe an analogous electron-hydrodynamic instability that produces similar running waves in Dirac materials such as graphene mono- and multilayers. It arises when carrier kinetics near charge neutrality make current dissipation strongly density-dependent. As the flow velocity $u$ exceeds a critical value, the system transitions to a state with coupled spatial and temporal oscillations. Experimentally, the instability should manifest as (i) a nonanalytic behavior characteristic of a second-order transition—an abrupt increase in time-averaged current—and (ii) narrow-band emission at the characteristic “washboard” frequency $f=u/\lambda$, where $\lambda$ is the modulation wavelength. This behavior parallels the AC and DC transport of sliding charge-density waves, but here it originates from a distinct, intrinsic mechanism unrelated to disorder. Estimates indicate that the emission frequency $f$, tunable by current, spans a broad range, highlighting Dirac bands as a promising platform for high-frequency electron-fluid dynamics.
@@
Figure 1: (a) Current dissipation in graphene bilayer (BLG) band vs. electron density. Dissipation is high in the Dirac plasma near charge neutrality point (CNP) but drops rapidly upon doping away from CNP. This behavior can drive the Turing–Kapitsa instability. Insets illustrate carrier population in BLG band for densities at CNP and away from it, respectively. (b) Instability growth rate ${\rm Im}\,\omega_{\pm}(k)$ for the modes $\omega_{+}(k)$ and $\omega_{-}(k)$, Eq. (18), representing density and velocity perturbations. The density mode satisfies $\omega_{+}(0)=0$ by particle-number conservation, becoming unstable above a critical flow velocity (Eq.(2)); the growth rate ${\rm Im}\,\omega_{+}(k)>0$ peaks at the wave numbers $\pm k_{*}$ marked by arrows. The velocity mode remains damped, ${\rm Im}\,\omega_{-}(k)<0$. Small initial perturbations grow into a periodic modulation with wavelength $\lambda\sim 2\pi/k_{*}$ propagating downstream, as shown in Fig. 2.
@@
There are several mechanisms reported in the literature that cause dissipation to decrease as carrier concentration increases. Here we focus on current dissipation in Dirac bands due to electron–hole (e–h) scattering at charge neutrality (CNP) [4, 5, 6, 7, 8, 9, 10, 11]. This dissipation channel is strong at CNP, where the Dirac plasma is e–h balanced, but it weakens rapidly once the carrier density is tuned away from neutrality. The resulting density-dependent dissipation plays a role similar to the width-dependent viscous drag in thin-film hydrodynamics: when the drift velocity exceeds a threshold, the flow becomes unstable and develops downstream-propagating waves. Our analysis predicts a critical drift velocity (see below):
$u_{c}=\frac{u_{0}}{R},\qquad R=\frac{n}{\gamma}\left|d\gamma/dn\right|,\qquad u_{0}=\sqrt{Un/m},$ (2)
where the dimensionless quantity $R$ characterizes how sensitively the dissipation rate $\gamma$ responds to changes in carrier density; $u_{0}$ is 2D plasmon velocity at density $n$. We emphasize that, because the dissipative regime of interest lies close to CNP, the relevant carrier densities are much smaller—i.e., much closer to neutrality—than those typically employed in studies of ballistic 2D plasmon propagation [38, 39, 40]. As illustrated in Fig. 1(a), the strong $T$-dependence of $\gamma$ leads to a pronounced $n$-dependence of $R$ at finite carrier densities near CNP. As discussed in [41], this causes the critical velocity $u_{c}$ to vary strongly with both $T$ and $n$.
The self-sustained modulation that appears for $u>u_{c}$ follows Turing’s wavelength-selection rule. Near threshold, the modulation wavelength is $\lambda=2\pi/k_{*}$, where $k_{*}$ is the wavenumber of the most unstable harmonic—i.e., the mode at which the instability growth rate is maximal (see Figs. 1(b) and 2). The transition is continuous and second-order (see Fig. 3). For realistic parameters, the resulting periodicity $\lambda$ can vary widely, from submicron to hundreds of microns [41].
@@
Figure 3: Signatures and observables for current-driven electronic Turing–Kapitsa instability. Plotted are the running waves modulation amplitude and time-averaged current near the instability onset, obtained numerically for a graphene bilayer model (see Eqs. (10), (11) and accompanying text). The red line shows the theoretical square-root behavior $\Delta n\propto(u_{d}-u_{d}^{*})^{1/2}$ expected on the general symmetry grounds [52] The sharp slope change in the time-averaged (DC) current–field dependence marks the instability threshold and the added current from the running waves. Insets (a) and (b) illustrate a typical modulation and the current–field relation.
@@
We are interested in transport in a Dirac band of a graphene bilayer, $\varepsilon_{\pm}(\mathbf{k})=\pm\tfrac{1}{2}\!\left(\sqrt{\gamma_{1}^{2}+4(\hbar vk)^{2}}-\gamma_{1}\right)$, where $v\approx 10^{6}\,$m/s and $\gamma_{1}=0.39\,$eV. For simplicity, we consider a one-species model representing one component of the bilayer bands—either electron or hole. The results below are largely insensitive to the exact form of $\epsilon(p)$ or to microscopic details of how momentum dissipation varies with carrier density. We model this by a dissipation rate that peaks in the region of width $w\sim k_{B}T$ and decreases with detuning from the CNP[10, 11]:
@@
$\left(\gamma(n)+\partial_{t}\right)$ ${\boldsymbol{p}}(\xi)+\partial_{{\boldsymbol{x}}}\Pi_{0}(\xi)=en\left({\boldsymbol{E}}-\partial_{{\boldsymbol{x}}}\Phi(\xi)\right)$
@@
The quantity $\Pi_{0}$ is the Fermi pressure of a flowing Fermi sea (FS). For Doppler-shifted Fermi distribution $f({\boldsymbol{p}})=\frac{1}{e^{\beta(\epsilon(p)-{\boldsymbol{u}}\cdot{\boldsymbol{p}}-\mu)}+1}$ describing electron gas of density $n$ flowing with the velocity $u$. For parabolic band, $\Pi_{0}$ is a sum of a static contribution and a kinetic energy term
@@
The transport equations in Eqs.(6) capture the essential physics underlying the current-driven Turing instability. A more realistic description, however, must explicitly include the interaction and viscous effects incorporated in the hydrodynamic stress density $\Pi$.
@@
Before analyzing amplification and instability in the strongly dissipative regime $\gamma\gg\omega$—the main focus of this work—it is useful to verify that, in the weak-dissipation limit, our results reproduce the known plasma mode dispersions, both with and without steady background flow. Setting $\gamma$, $\Lambda$, $D_{n}$ and $D_{p}$ to zero gives
@@
Next, we return to the analysis of dissipative transport and estimate maximally unstable wavelength and instability threshold. It is instructive to focus on the strongly dissipative regime, where the momentum relaxation rate $\gamma$ greatly exceeds scales such as $\omega p$ and $k\Pi$. For the parabolic band model the general expression for $\omega_{\pm}$ given in Eqs. (18) and (19) simplifies considerably. Substituting the derivatives $\Pi_{n}$ and $\Pi_{p}$ in Eq. (19) yields
$\Delta=\gamma^{2}+ik\frac{4p}{m}\frac{d\gamma}{dn}-k^{2}\frac{4Un}{m}$ (22)
An analytic estimate of the most unstable wavelength $k_{*}$ can be obtained by examining the instability growth rate ${\rm Im}\,\omega_{+}(k)$ vs. $k$ in the limits of small and large $k$, and matching the two asymptotic behaviors in the intermediate (“borderline”) region. For clarity, we focus on the case of small diffusion coefficients $D_{n}$ and $D_{p}$, setting $D_{n}=D_{p}=0$. In the large-$k$ limit ($k\gg k_{*}$), the growth rate approaches a $k$-independent value:
${\rm Im}\,\omega_{+}(k)_{k\gg k_{*}}=-\frac{1}{2}\gamma+\frac{s}{2u_{0}}\frac{pd\gamma}{mdn},\quad s={\rm sgn}\,\left(p\frac{d\gamma}{dn}\right).$ (23)
This predicts the instability threshold at
$p\frac{d\gamma}{dn}=s\gamma mu_{0}$ (24)
with an unstable phase for $p\frac{d\gamma}{dn}s>\gamma mu_{0}$ and a stable phase for $p\frac{d\gamma}{dn}<\gamma mu_{0}$. Namely, the instability occurs when the flow velocity $u=p/mn$ exceeds the critical value $u_{c}$ given in Eq.(2). If $R$ is order-one, the flow velocity at which instability occurs is comparable to the plasmonic velocity $u_{0}=\sqrt{Un/m}$.
Next, to complete the estimate of the most unstable wavelength we consider small $k$, where Taylor expansion of ${\rm Im}\,\omega_{+}(k)$ to leading order in $k$ gives
${\rm Im}\,\omega_{+}(k)_{k\ll k_{*}}=k^{2}\left[\frac{p^{2}\left(\frac{d\gamma}{dn}\right)^{2}}{m^{2}\gamma^{3}}-\frac{u_{0}^{2}}{\gamma}\right]+O(k^{4})$ (25)
For $k\approx k_{*}$, these values should match those in Eq.(23), obtained at $k\gg k_{*}$. Writing this condition as
$k_{*}^{2}\left[\frac{p^{2}\left(\frac{d\gamma}{dn}\right)^{2}}{m^{2}\gamma^{3}}-\frac{u_{0}^{2}}{\gamma}\right]=\frac{p\frac{d\gamma}{dn}}{2mu_{0}}-\frac{1}{2}\gamma$ (26)
and noting that both the left- and the right-hand side vanish at the transition but remain nonzero away from transition, we obtain [51]:
@@
Notably, the maximally unstable wavelength, $\lambda_{*}=2\pi/k_{*}$, remains finite at the transition. This behavior is consistent with our simulation results and mirrors the finite modulation wavelength observed at the onset of the hydrodynamic Kapitsa instability[34, 35]. Density-dependent $\gamma$ and $u_{0}$ lead to tunable drift velocity $u_{d}$ and wavelength spanning a range of values, from submicron to hundreds of microns. Accordingly, the narrow-band radiation frequency, given by the washboard formula $f=u_{d}/\lambda$, for realistic system parameters can vary from sub-GHz to hundreds of GHz. For estimates and plots of critical velocity and modulation wavelength see [41].
In summary, we identify an electron–hydrodynamic instability in Dirac materials that is closely analogous to the Turing-Kapitsa instability generating trains of roll waves in viscous films. In electron systems, strong density dependence of current dissipation near charge neutrality makes the flow susceptible to modulation, triggering a transition at a critical drift velocity $u_{c}$, beyond which the system develops self-sustained running waves. This instability produces both a sharp, second-order–like change in the time-averaged current and narrow-band emission at the washboard frequency $f=u/\lambda$. Estimated frequencies span a wide, current-tunable range extending to hundreds of GHz, highlighting Dirac materials as a promising platform for high-frequency electron-fluid phenomena.
This work greatly benefited from discussions with Ray Ashoori, Serguei Brazovskii, Andre Geim, Frank Koppens, Brian Skinner and Eli Zeldov.

### v2

@@ Abstract
We predict hydrodynamic Turing instability of current-carrying Dirac electron fluids that drives spontaneous self-oscillatory transport. The instability arises near charge neutrality, where carrier kinetics make current dissipation strongly density dependent. Above a critical drift velocity, a uniform electronic flow becomes unstable and undergoes a dynamical transition to a state with coupled spatial modulation and temporal oscillations—an electronic analogue of Kapitsa roll waves in viscous films. The transition exhibits two clear signatures: a nonanalytic, second-order–like onset in the time-averaged current and narrow-band electromagnetic emission at a tunable washboard frequency $f=u/\lambda$. Although reminiscent of sliding charge-density waves, the mechanism is intrinsic and disorder independent. Owing to the small effective mass of Dirac carriers, hydrodynamic time scales translate into emission frequencies in the tens to hundreds of gigahertz range, establishing Dirac materials as a platform for high-frequency self-oscillatory electron hydrodynamics.
@@
Figure 1: (a) Current dissipation in graphene bilayer (BLG) band vs. electron density. Dissipation is high in the Dirac plasma near charge neutrality point (CNP) but drops rapidly upon doping away from CNP. Density-dependent dissipation drives the Turing–Kapitsa instability. Insets illustrate carrier population in BLG band for densities at CNP and away from it, respectively. (b) Instability growth rate ${\rm Im}\,\omega_{\pm}(k)$ for the modes $\omega_{+}(k)$ and $\omega_{-}(k)$, Eq. (19), representing density and velocity perturbations. The density mode satisfies $\omega_{+}(0)=0$ by particle-number conservation, becoming unstable above a critical flow velocity (Eq.(2)); the growth rate ${\rm Im}\,\omega_{+}(k)>0$ peaks at the wave numbers $\pm k_{*}$ marked by arrows. The velocity mode remains damped, ${\rm Im}\,\omega_{-}(k)<0$. Small initial perturbations grow into a periodic modulation with wavelength $\lambda\sim 2\pi/k_{*}$ propagating downstream, as shown in Fig. 2.
@@
Several effects reported in the literature can cause dissipation to decrease as carrier concentration increases. Here we focus on quasiparticle scattering at charge neutrality (CNP) [4, 5, 6, 7, 8, 9, 10, 11]. This dissipation channel is strong at CNP, where the Dirac plasma is e–h balanced, but it weakens rapidly once the carrier density is tuned away from neutrality. The resulting density-dependent dissipation plays a role similar to the width-dependent viscous drag in thin-film hydrodynamics: when the drift velocity exceeds a threshold, the flow becomes unstable and develops downstream-propagating waves.
Our analysis predicts a critical drift velocity:
$u_{c}=\frac{u_{0}}{R},\qquad R=\frac{n}{\gamma}\left|\frac{d\gamma}{dn}\right|,\qquad u_{0}=\sqrt{\frac{Un}{m}},$ (2)
where the dimensionless quantity $R$ measures how sensitively the dissipation rate $\gamma$ responds to changes in carrier density, while $u_{0}$ denotes the 2D plasmon velocity at density $n$. Because the dissipative regime of interest lies close to the charge neutrality point (CNP), the relevant carrier densities are much smaller—i.e., much closer to neutrality—than those typically used in studies of ballistic 2D plasmon propagation [29, 30, 31]. As shown in Fig. 1(a), the strong temperature dependence of $\gamma$ produces a pronounced density dependence of $R$ near the CNP. As a result, the critical velocity $u_{c}$ varies strongly with both temperature and carrier density [32].
For $u>u_{c}$, the system develops a self-sustained modulation that follows Turing’s wavelength-selection rule. Close to threshold, the modulation wavelength is $\lambda=2\pi/k_{*}$, where $k_{*}$ is the most unstable mode, at which the instability growth rate is maximal (see Figs. 1(b) and 2). The transition is continuous and second-order (Fig. 3), and the modulation wavelength remains finite at the threshold, as in Brazovskii and Swift-Hohenberg theories of convective instabilities[33, 34], without showing any divergence. For realistic parameters, $\lambda$ spans a wide range, from submicron to hundreds of microns [32].
@@
Figure 3: Signatures and observables for current-driven electronic Turing–Kapitsa instability. Plotted are the running waves modulation amplitude and time-averaged current near the instability onset, obtained numerically for a graphene bilayer model (see Eqs. (11), (12) and accompanying text). The red line shows the theoretical square-root behavior $\Delta n\propto(u_{d}-u_{d}^{*})^{1/2}$ expected on the general symmetry grounds [45]. The sharp slope change in the time-averaged (DC) current–field dependence marks the instability threshold and the added current from the running waves. Insets (a) and (b) illustrate a typical modulation and the current–field relation.
@@
We are interested in transport in a Dirac band of a graphene bilayer, $\varepsilon_{\pm}({\boldsymbol{k}})=\pm\tfrac{1}{2}\!\left(\sqrt{\gamma_{1}^{2}+4(\hbar vk)^{2}}-\gamma_{1}\right)$, where $v\approx 10^{6}\,$m/s and $\gamma_{1}=0.39\,$eV. For simplicity, we consider a one-species model representing one component of the bilayer bands—either electron or hole. The results below are largely insensitive to the exact form of $\epsilon(p)$ or to microscopic details of how momentum dissipation varies with carrier density. We model this by a dissipation rate that peaks in the region of width $w\sim k_{B}T$ and decreases with detuning from the CNP[10, 11]:
@@
$\left(\gamma(n)+\partial_{t}\right)$ ${\boldsymbol{p}}(\xi)+\partial_{{\boldsymbol{x}}}\Pi_{0}(\xi)=en\left({\boldsymbol{E}}-\partial_{{\boldsymbol{x}}}\Phi(\xi)\right)$ (6)
@@
The quantity $\Pi_{0}$ is the Fermi pressure of a flowing Fermi sea described by Doppler-shifted Fermi distribution $f({\boldsymbol{p}})=\frac{1}{e^{\beta(\epsilon(p)-{\boldsymbol{u}}\cdot{\boldsymbol{p}}-\mu)}+1}$. For electron gas of density $n$ flowing with the velocity $u$, for parabolic band, $\Pi_{0}$ is a sum of a static contribution and a kinetic energy term
@@
The transport equations (6) and (7) capture the essential physics underlying the current-driven Turing instability. A more realistic description, however, must explicitly include the interaction and viscous effects incorporated in the hydrodynamic stress density $\Pi$.
@@
Figure 4: Modulation wavelength $\lambda=2\pi/k_{*}$, where $\pm k_{*}$ (blue arrows) maximizes the instability growth rate $\mathrm{Im}\,\omega_{+}(k)$, is governed by the behavior of $\mathrm{Im}\,\omega_{+}(k)$ at small and large $k$. For small $k$, the growth rate (red parabola) behaves as $\alpha k^{2}$, with $\alpha>0$ for $u>u_{c}$ and $\alpha<0$ for $u<u_{c}$. For large $k$, it saturates to a $k$-independent value (horizontal red line), again positive for $u>u_{c}$ and negative for $u<u_{c}$. Matching these small- and large-$k$ asymptotics yields Eq. (24), and hence Eq. (26). The instability growth rate is shown for $u=0.55>u_{c}$ and $u=0.2<u_{c}$ (blue and pink curves); solid and dashed lines correspond to $D_{n,p}=0.1$ and $D_{n,p}=0$.
Prior to analyzing the dissipative regime $\gamma\gg\omega$—the main focus of this work—we verify that, in the weak-dissipation limit, our results reproduce the known plasma mode dispersions, both with and without steady background flow. Setting $\gamma$, $\Lambda$, $D_{n}$ and $D_{p}$ to zero gives
@@
Next, we estimate maximally unstable wavelength and instability threshold for the dissipative regime, where the relaxation rate $\gamma$ greatly exceeds scales such as $\omega p$ and $k\Pi$. For a parabolic band model the general expression for $\omega_{\pm}$ given in Eqs. (19) and (20) simplifies considerably. Substituting the derivatives $\Pi_{n}$ and $\Pi_{p}$ in Eq. (20) yields
$\Delta=\gamma^{2}+ik\frac{4p}{m}\frac{d\gamma}{dn}-4k^{2}u_{0}^{2}$ (23)
An analytic estimate of the most unstable wavenumber $k_{\ast}$ can be obtained as illustrated in Fig. 4. We compare the instability growth rate ${\rm Im}\,\omega_{+}(k)$ in the small- and large-$k$ limits and match these asymptotic behaviors in an intermediate (“borderline”) region. For clarity, we focus on the case of small $D_{n}$ and $D_{p}$. In the large- and small-$k$ limits, the growth rate behaves as
${\rm Im}\,\omega_{+}(k)=\begin{cases}-\frac{1}{2}\gamma+\frac{\left|p\frac{d\gamma}{dn}\right|}{2mu_{0}},&k\gg k_{*},\\ k^{2}\frac{\left[p^{2}\left(\frac{d\gamma}{dn}\right)^{2}-m^{2}\gamma^{2}u_{0}^{2}\right]}{m^{2}\gamma^{3}}&k\ll k_{*}.\end{cases}$ (24)
The sign change predicts instability threshold $u=u_{c}$ with the unstable and stable phases for $\left|p\frac{d\gamma}{dn}\right|-\gamma mu_{0}$ positive and negative, respectively. The sign change occurs simultaneously for $k\gg k_{*}$ and $k\ll k_{*}$. This gives critical velocity $u_{c}$ given in Eq.(2). If $R\sim 1$, the critical velocity is comparable to the plasmonic velocity $u_{0}$ in a dissipation-free system at density close to CNP.
An estimate of the most unstable mode $k_{\star}$ follows from matching the small- and large-$k$ behavior Eq.(24):
$\frac{k_{*}^{2}}{m^{2}\gamma^{3}}\left[p^{2}\left(\frac{d\gamma}{dn}\right)^{2}-m^{2}\gamma^{2}u_{0}^{2}\right]=\frac{\left|p\frac{d\gamma}{dn}\right|}{2mu_{0}}-\frac{1}{2}\gamma.$ (25)
Solving for $k_{*}$ and noting that both sides vanish at the transition but remain nonzero away from transition, in the limit $\left|p\frac{d\gamma}{dn}\right|\to mu_{0}\gamma$ we obtain [44]:
@@
Notably, the maximally unstable wavelength $\lambda_{\ast}=2\pi/k_{\ast}$ remains finite at the transition, in agreement with our simulations and in analogy with the finite modulation wavelength at the onset of the hydrodynamic Kapitsa instability.[25, 26] Density-dependent $\gamma$ and $u_{0}$ yield a tunable critical velocity $u_{c}$ and wavelength ranging from submicron to hundreds of microns. The resulting narrow-band radiation frequency, $f=u_{d}/\lambda$, can thus span from sub-GHz to hundreds of GHz for realistic parameters; see [32] for estimates and plots of the critical velocity and modulation wavelength.
In summary, we identify an electron–hydrodynamic instability in Dirac materials closely analogous to the Turing–Kapitsa instability that generates roll-wave trains in viscous films. Density dependence of current dissipation near charge neutrality makes the flow unstable above a critical drift velocity $u_{c}$, resulting in self-sustained running waves. This leads to a sharp, second-order–like change in the time-averaged current and narrow-band emission at the washboard frequency $f=u/\lambda$, with a current-tunable range extending to hundreds of GHz, highlighting Dirac materials as a promising platform for high-frequency electron-fluid phenomena.
This work greatly benefited from discussions with Ray Ashoori, Serguei Brazovskii, Andre Geim, Frank Koppens, Leo Radzihovsky, Brian Skinner, and Eli Zeldov.